### v1

@@ IV-A Inter-Annotator Agreement
To assess the reliability of human and LLM annotations, a subset of 339 instances from the test set, balanced by labels and languages, was annotated by a second annotator. The nominal Krippendorff’s alpha [43] between the two annotators reached 0.728, surpassing the acceptable threshold of 0.667, as defined by Krippendorff [43]. This moderate level of agreement indicates that the annotation campaign was successful while also highlighting the challenging nature of this task. As shown in Table II, the inter-annotator agreement between each of the human annotators and the LLM model was on par with the agreement between the two human annotators. This finding addresses Research Question 1 (RQ1), demonstrating that the LLM model is as capable of data annotation for this task as the human annotators.
@@
We also assess the agreement between human and LLM annotators at the label level. Fig. IV-A illustrates the substantial variation in inter-annotator agreement across topic labels, highlighting labels that are less clearly defined, which leads to confusion among annotators. The labels sport and weather are well comprehended by both human and LLM annotators, achieving high inter-annotator agreement above 0.90 in terms of Krippendorff’s alpha. Labels religion, lifestyle and leisure, and society present challenges for the LLM, with human annotators demonstrating significantly higher agreement. The label economy, business and finance is particularly problematic for both humans and the LLM, with an average inter-annotator agreement of 0.43.
@@ IV-B Intra-Annotator Agreement
To further assess the reliability of the annotations provided by humans and LLMs, we instruct both a human annotator and the GPT-4o model to re-annotate the subset that was previously used to calculate the inter-annotator agreement. The annotators are given the same instructions as in the initial annotation process, and the labels previously assigned to the texts are hidden from them. Intra-annotator agreement is evaluated using Krippendorff’s alpha measuring consistency between two annotation runs conducted by the same annotator. Table III presents the intra-annotator agreement for the human and LLM annotators. The human annotator achieved a relatively high Krippendorff’s alpha of approximately 0.80, indicating a reliable level of self-agreement. The GPT-4o model demonstrated even greater consistency with Krippendorff’s alpha of 0.93, highlighting the reliability of large language models in producing consistent responses. However, while such consistency is also advantageous, more diverse responses, similar to those obtained from the human annotator, can be beneficial for various research directions, including learning from disagreement [44].
@@ V Student Model Fine-Tuning Experiments
In this section, we describe the experiments involving the fine-tuning of student BERT-like models on training data that were automatically annotated by the teacher GPT model. In Section V-A, we present the student model is used in the experiments that examine (1) the impact of training data size on classification performance (Section V-B), (2) the model’s label-level performance (Section V-C), and (3) the model’s monolingual, multilingual, and cross-lingual capabilities (Section V-D).
@@ V-C Label-Level Performance
The best-performing version of the XLM-RoBERTa student model, namely, a version trained on 15,000 instances, has been made available in the Hugging Face repository (accessible at https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier). This model attains a macro-F1 score of 0.746 and micro-F1 score of 0.734. Further insights into the performance of the model at the label level can be obtained from the confusion matrix depicted in Fig. V-B, which displays the true versus predicted classifications.
The labels that are most accurately predicted by the model are weather, sport, and religion, with label-level F1 scores above 0.89. These findings are consistent with the label-level inter-annotator agreement shown in Fig. IV-A (see Section IV-A), which identified these labels as those with the highest consensus among human annotators. Interestingly, despite the weather label being represented by less than 1% of the instances of the training dataset, this did not negatively impact the performance of the model. The model demonstrated a near-perfect classification accuracy for this label, achieving an F1 score of 0.94.
In contrast, the most challenging labels for the XLM-RoBERTa student model are lifestyle and leisure, human interest, and economy, business and finance, as evidenced by label-level F1 scores ranging from 0.59 to 0.62. This is not surprising, given that these categories have also been identified as challenging for both human and LLM annotators. As shown in Fig. V-B, the label economy, business and finance is frequently misclassified as science and technology when the texts pertain to the sales of technology products; or as environment in instances where the texts address electricity infrastructure. The label human interest is often mistaken for arts, culture, entertainment and media, because both categories can encompass content related to celebrities. The former is typically associated with tabloid-style presentations of celebrity lives, whereas the latter is intended to cover more serious topics. Similarly, the label lifestyle and leisure frequently overlaps with both human interest and arts, culture, entertainment and media, as it encompasses individuals’ recreational activities, which may intersect with topics represented by the other two labels. These overlaps illustrate the complexities inherent in IPTC topic classification, which pose challenges to both manual annotation and automatic classification. These challenges arise primarily from the broad nature of top-level labels, which often leads to intersections among categories. A potential solution to these issues is to conduct classification with labels from the lower levels of the Media Topic schema, which are more specific but also more challenging due to their high number.
@@ V-D Monolingual, Multilingual and Cross-Lingual Performance
The models are evaluated on the manually-annotated test dataset, split into target language subsets. Each monolingual test subset comprises approximately 300 instances. The reported scores are averaged across the three iterations of fine-tuning and evaluation.
@@ VI Conclusion
The IPTC topic training dataset is developed by extracting news articles from web corpora in four languages (Slovenian, Croatian, Greek, and Catalan) based on automatic genre identification using the X-GENRE classifier [1] and automatic topic annotation using the GPT model as a teacher model. The EMMediaTopic dataset [41] has been made freely available in the CLARIN.SI repository (http://hdl.handle.net/11356/1991).
@@
In this study, we employ only the 17 top-level categories of the IPTC Media Topic hierarchical schema. The analysis of the agreement between human annotators and the LLM as data annotator for each category, as well as the confusion matrix for the fine-tuned model, highlighted challenges associated with the ambiguity of the top-level labels, which, in certain instances, are not clearly delineated. In future work, we plan to extend our experiments to include more fine-grained categories from the lower levels of the IPTC schema, which are more useful for some users and news media providers. We consider following the hierarchical approach proposed by [20] and [19], where the top label is assigned to the text first with the classifier presented in this paper, and then based on the predicted label, a more specialized lower-layer classifier is selected to make deeper level-two predictions.

### v2

@@ IV-A Inter-Annotator Agreement
To assess the reliability of human and LLM annotations, a subset of 339 instances from the test set, balanced by labels and languages, was annotated by a second annotator. Human annotators received the same guidelines, provided in Appendix A, and participated in a preliminary test annotation which was followed by a discussion to resolve the disagreements.
The inter-annotator agreement is evaluated using Krippendorff’s alpha metric [44], a statistical measure frequently employed in content analysis, social sciences, and machine learning to determine the reliability of multiple human annotators in data classification. The metric accounts for the likelihood of agreement occurring by chance, with a value of 1 indicating perfect agreement and a value of 0 indicating agreement equal to chance [45]. Theoretically, reliable inter-annotator agreement is represented by alpha values of 0.8 or higher, whereas an alpha of 0.667 is considered the threshold for acceptable annotation reliability [44]. Nonetheless, this threshold value is frequently not achieved in manual annotation campaigns for text classification tasks. This shortcoming can be attributed to the high subjectivity and complexity of annotation tasks, which often involve cases where characteristics of multiple labels are present [46]. This problem has also been demonstrated in tasks similar to topic classification, such as automatic genre identification [47], sentiment annotation [48], [49, 50], and detection of hate speech [51, 52].
In our experiments, the nominal Krippendorff’s alpha between the two annotators reached 0.728. This outcome is deemed satisfactory as it exceeds the threshold for acceptable reliability. Furthermore, this level of inter-annotator agreement is comparable to or exceeds the agreement observed in similar text-level classification tasks. At the same time, the moderate level of agreement highlights the challenging nature of this task, as multiple topic labels may be discussed in certain news articles.
As shown in Table II, the inter-annotator agreement between each of the human annotators and the LLM model was on par with the agreement between the two human annotators. This finding addresses Research Question 1 (RQ1), demonstrating that the LLM model is as capable of data annotation for this task as human annotators.
@@
We also assess the agreement between the human and LLM annotators at the label level. Fig. IV-A illustrates the substantial variation in inter-annotator agreement across topic labels, highlighting labels that are less clearly defined, which leads to confusion among annotators. The labels sport and weather are well comprehended by both human and LLM annotators, achieving high inter-annotator agreement above 0.90 in terms of Krippendorff’s alpha. Labels religion, lifestyle and leisure, and society present challenges for the LLM, with human annotators demonstrating significantly higher agreement. The label economy, business and finance is particularly problematic for both humans and the LLM, with an average inter-annotator agreement of 0.43.
@@ IV-B Intra-Annotator Agreement
To further assess the reliability of the annotations provided by humans and LLMs, we instruct both a human annotator and the GPT-4o model to re-annotate the subset that was previously used to calculate the inter-annotator agreement. Annotators are given the same instructions as in the initial annotation process, and the labels previously assigned to the texts are hidden from them. Intra-annotator agreement is evaluated using Krippendorff’s alpha measuring consistency between two annotation runs conducted by the same annotator. Table III presents the intra-annotator agreement for the human and LLM annotators. The human annotator achieved a relatively high Krippendorff’s alpha of approximately 0.80, indicating a reliable level of self-agreement. The GPT-4o model demonstrated even greater consistency with Krippendorff’s alpha of 0.93, highlighting the reliability of large language models in producing consistent responses. However, while such consistency is advantageous, more diverse responses, similar to those obtained from the human annotator, can be beneficial for various research directions, including learning from disagreement [53].
@@ V Student Model Fine-Tuning Experiments
In this section, we describe the experiments that involve fine-tuning of student BERT-like models on training data that were automatically annotated by the teacher GPT model. In Section V-A, we present the student model which is used in the experiments that examine (1) the impact of training data size on classification performance (Section V-B), (2) the model’s label-level performance (Section V-C), and (3) the model’s monolingual, multilingual, and cross-lingual capabilities (Section V-D).
@@ V-C Label-Level Performance
The best-performing version of the XLM-RoBERTa student model, namely, a version trained on 15,000 instances, has been made available in the Hugging Face repository (accessible at https://huggingface.co/classla/multilingual-IPTC-news-topic-classifier). This model attains a macro-F1 score of 0.746 and micro-F1 score of 0.734. Further insights into the performance of the model at the label level can be obtained from the confusion matrix depicted in Fig. V-B, which displays the true versus predicted labels.
The labels that are most accurately predicted by the model are weather, sport, and religion, with label-level F1 scores above 0.89. These findings are consistent with the label-level inter-annotator agreement shown in Fig. IV-A (see Section IV-A), which identified these labels as those with the highest consensus among human annotators. Interestingly, despite the weather label being represented by less than 1% of the instances of the training dataset, this did not negatively impact the performance of the model. The model demonstrated a near-perfect classification accuracy for this label, achieving an F1 score of 0.94. This result confirms the usefulness of the classifier also for the labels that are less represented in the training dataset.
In contrast, the most challenging labels for the XLM-RoBERTa student model are lifestyle and leisure, human interest, and economy, business and finance, as evidenced by label-level F1 scores ranging from 0.59 to 0.62. This is not surprising, given that these categories have also been identified as challenging for both human and LLM annotators. As shown in Fig. V-B, the label economy, business and finance is frequently misclassified as science and technology when the texts pertain to the sales of technology products; or as environment in instances where the texts address electricity infrastructure. The label human interest is often mistaken for arts, culture, entertainment and media, because both categories encompass content related to celebrities. The former is typically associated with tabloid-style presentations of celebrity lives, whereas the latter is intended to cover more serious topics. Similarly, the label lifestyle and leisure frequently overlaps with both human interest and arts, culture, entertainment and media, as it encompasses individuals’ recreational activities, which may intersect with topics represented by the other two labels. These overlaps illustrate the complexities inherent in IPTC topic classification, which pose challenges to both manual annotation and automatic classification. These challenges arise primarily from the broad nature of top-level labels, which often leads to intersections among categories. A potential solution to these problems is to conduct classification with labels from the lower levels of the Media Topic schema, which are more specific but also more challenging due to their high number.
@@ V-D Monolingual, Multilingual and Cross-Lingual Performance
The models are evaluated on the manually annotated test dataset, split into target language subsets. Each monolingual test subset comprises approximately 300 instances. The reported scores are averaged across the three iterations of fine-tuning and evaluation.
@@ VI Conclusion
The IPTC Media Topic training dataset is developed by extracting news articles from web corpora in four languages (Slovenian, Croatian, Greek, and Catalan) based on automatic genre identification using the X-GENRE classifier [1] and automatic topic annotation using the GPT model as a teacher model. The EMMediaTopic dataset [42] has been made freely available in the CLARIN.SI repository (http://hdl.handle.net/11356/1991).
@@
In this study, we employ only the 17 top-level categories of the IPTC Media Topic hierarchical schema. The analysis of the agreement between human annotators and the LLM annotator for each category, as well as the confusion matrix for the fine-tuned model, highlighted challenges associated with the ambiguity of certain top-level labels. In future work, we plan to extend our experiments to include more fine-grained categories from the lower levels of the IPTC schema, which are more useful for some users and news media providers. We consider following the hierarchical approach proposed by [20] and [19], where the top label is assigned to the text first with the classifier presented in this paper, and then based on the predicted label, a more specialized lower-layer classifier is selected to make deeper level-two predictions. Moreover, our current experiments involve single-label classification, which does not account for the multifaceted nature of some texts in which multiple top-level topics could be useful to the end user. To address this, we plan to move towards a multi-label classification setting based on the information on the confidence of the student model.
@@
Furthermore, future work could include a more in-depth exploration of potential biases that might emerge in the fine-tuned student model. The biases may be introduced from the original sources, since the training data is sourced from a large collection of online news articles, as well as from the large language model used for annotating these training data. Given that the news topic categories include sensitive subjects such as conflict, war and peace, religion, and society, it is important to consider the potential biases originating from both the training data and the large language model while deploying the automatic news classification system in real-world applications.
VII Limitations
While our LLM teacher-student methodology provides promising results, it is important to acknowledge that the success of this approach is fundamentally dependent on the credibility of the teacher model, as its performance represents the upper limit of what can be achieved with the framework. In our study, the GPT model demonstrated strong performance in the news topic classification task, achieving high macro-F1 and micro-F1 scores and exhibiting annotation reliability comparable to that of human annotators. These findings highlight the feasibility of the framework for this specific task, providing high-quality annotations at minimal cost. However, this level of performance may not necessarily be replicable across other natural language processing tasks or languages, where the conventional method of training on manually annotated data may still be preferable. Therefore, it is crucial that researchers assess the teacher model on manually annotated test datasets to ensure its reliability and comparability to human annotators prior to applying the LLM teacher-student approach to their use cases.